OpenStudy (anonymous):

You decide to save money to pay the fees for your hobby classes and save $22 in the first week. If you save $3 more each week than in the preceding week, how much money will you have by the end of the 10th week? A.) $340 B.) $355 C.) $381 D.) $407 E.) $435

OpenStudy (anonymous):

22 + [22 + 3(1)] + [22 + 3(2)] + . . . + [22 + 3(9)] = (10)(22) + 3(1 + 2 + 3 + . . . + 9) = (10)(22) + 3(9)(10)/2 = ?

OpenStudy (anonymous):

355?

OpenStudy (anonymous):

Yep! You got it!

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

uw! Good luck in all of your studies and thx for the recognition! @SlyWolf0122

OpenStudy (anonymous):

Hey @tcarroll010 I just want to say that your helping @SlyWolf0122 cheat on a big exam he has to take on a web page named ple.platoweb.com because he failed a semester of algbra 1.And @SlyWolf0122 THIS WEB PAGE IS FOR HELP, NOT 4 ANSWERS ON A TEST.